# Scalability Challenges in Variational Quantum Optimization under Stochastic Noise

### v1

@@ Abstract
With rapid advances in quantum hardware, a key question is whether quantum devices without full error correction can outperform classical computers on practically relevant problems. Variational Quantum Algorithms (VQAs) have gained significant attention as promising candidates in this pursuit, particularly for combinatorial optimization problems. While reports of their challenges and limitations continue to accumulate, many studies still convey optimism based on small-scale toy problems and idealized testing setups. However, doubts remain about the scalability of VQAs and hence their viability for real-world problems. We systematically investigate this scaling behavior by analyzing how state-of-the-art classical optimizers minimize well-behaved quantum loss functions for random QUBO problem instances. Special emphasis is placed on how these algorithms handle uncertainties, modeled as effective Gaussian noise. Our findings reveal that the critical noise threshold, beyond which classical optimizers fail to find optimal or near-optimal solutions, decreases rapidly with system size. This effect exceeds what can be attributed to barren plateaus, indicating more fundamental limitations inherent to the hybrid paradigm of VQAs. When translating this threshold to the finite sampling error permissible in a fault-tolerant scenario, the required number of measurement shots becomes prohibitively high, even for medium-sized problems. These results raise serious doubts about the feasibility of achieving a practical quantum advantage in optimization using VQAs. Instead, they highlight the urgent need to explore fundamentally different algorithms and problem domains that can overcome the reported scalability challenges.
@@ I Introduction
While quantum hardware has advanced rapidly [1, 2, 3], scalable and fault-tolerant quantum computing may still be a long way off. John Preskill [4] coined the term NISQ (Noisy Intermediate-Scale Quantum) to describe a generation of quantum devices, which operate with limited resources and significant noise levels. A key question in the NISQ era is whether these early-stage devices can achieve a practically relevant quantum advantage. In this context, variational quantum algorithms (VQAs) have emerged as potentially promising candidates for realizing near-term quantum advantage in a wide range of applications [5, 6, 7]. One of the most anticipated use cases, particularly in industry, is combinatorial optimization, where current quantum approaches mainly rely on approximation algorithms and heuristics such as VQAs [8].
@@
VQAs are hybrid schemes that leverage quantum computers to evaluate a loss function while relying on classical optimization methods for loss minimization. They employ shallow, parameterized quantum circuits (ansatzes), making them well-suited to the resource constraints of NISQ devices. However, a growing body of research highlights significant challenges and limitations of VQAs. One major issue is the barren plateau phenomenon [9, 10], where the loss values and gradients vanish exponentially with system size in most of the proposed quantum ansatzes, making the parameter optimization intractable. Conversely, structural properties leading to the provable absence of barren plateaus might enable efficient classical simulations [11]. Additionally, the classical optimization problem within VQAs has been proven to be NP-hard [12], with loss landscapes plagued by numerous suboptimal local minima [13], complicating the convergence to high-quality solutions. These challenges are amplified by noise, which not only leads to unfavorable resource scaling and unpredictable behavior but may also eliminate any potential quantum advantage altogether [14, 15, 16, 17, 18, 19, 20]. Crucially, these limitations persist regardless of whether the effective noise comes from hardware imperfections or finite sampling errors inherent to quantum measurement. Consequently, despite ongoing efforts and widespread optimism, no practically meaningful quantum advantage has been demonstrated for VQAs on near-term devices to date.
@@
Thus, the critical question remains: can variational quantum optimization become practical in the presence of unavoidable stochastic errors? While some studies directly address this question [21, 16, 17, 18], most research focuses on developing new heuristic approaches, typically tested on small-scale and idealized setups. Such benchmarks often yield promising insight but rarely capture how well algorithms scale under realistic noise conditions and increasing problem sizes, potentially painting an overly optimistic picture of their viability. More importantly, many studies overlook the role of the classical optimization subroutine – the central computational bottleneck of VQAs – which may ultimately determine whether a problem remains solvable as system size grows. Closing this gap is essential for distinguishing genuinely promising advancements from appealing but unproven claims in quantum optimization and related disciplines.
@@
This work critically examines the viability of VQAs for solving classical optimization problems in a Quadratic Unconstrained Binary Optimization (QUBO) formulation, with a specific focus on the classical optimization process as the key driver of performance in these hybrid algorithms. By systematically analyzing the ability of state-of-the-art classical optimizers to minimize quantum loss functions under increasing levels of noise, we provide a comprehensive and generalizable assessment of the practical limits of variational quantum optimization. The key contributions of this work are:
@@
    General Framework for Evaluation: We establish a systematic methodology for assessing the performance of quantum optimization algorithms, providing a robust foundation for analyzing their potential and limitations under realistic conditions.
@@
    Scaling Analysis of Classical Optimizers: To the best of our knowledge, this study presents the first in-depth scaling analysis of classical optimizer performance on noisy variational quantum loss functions for random QUBO problem instances. By systematically testing across two key dimensions – noise level $\sigma$ and system size $n$ – we identify a clear threshold $\sigma^{*}(n)$, beyond which optimization problems become unsolvable. Notably, the scaling analysis allows extrapolating to regions not (yet) accessible through numerical testing.
@@
    Implications for Practical Quantum Advantage: Our findings offer quantitative insights into how stochastic errors, including those arising solely from finite measurement sampling, fundamentally limit the feasibility of VQAs. The results strongly challenge the feasibility of achieving quantum advantage in variational quantum optimization under realistic conditions.
Through rigorous empirical validation, this work provides experimental evidence that highlights the practical limitations of VQAs under noise, offering a much-needed benchmark for evaluating their scalability and true potential. By fostering a more realistic understanding of quantum optimization on near-term devices and beyond, these insights help to guide future research towards more promising algorithms and applications in quantum computing.
II Theoretical Background
@@
II.1 Quadratic Unconstrained Binary Optimization Problems
@@
II.2 Variational Quantum Optimization
@@
Fig. 1 illustrates representative ansatz circuits for both algorithm classes, which are also used in the experiments presented in Section III.2. For VQE, a typical example is a hardware-efficient ansatz, consisting of alternating layers of single-qubit rotations and entangling operations, followed by a global measurement. QAOA, in contrast, follows a problem-inspired approach, implementing a trotterized version of adiabatic quantum optimization [27]. It consists of alternating applications of a shift operator ${\mathbf{U}_{P}(\gamma)=e^{-i\gamma\mathbf{H}_{P}}}$ based on the problem Hamiltonian $\mathbf{H}_{P}$, and a mixing operator ${\mathbf{U}_{M}(\beta)=e^{-i\beta\mathbf{H}_{M}}}$, where the default choice is a sum of Pauli-$\mathbf{X}$ matrices on each qubit, i.e., ${\mathbf{H}_{M}=\sum_{i=1}^{n}\mathbf{X}_{i}}$. The resulting QAOA ansatz is composed of $p$ layers, parameterized by $2p$ angles, with the unitary evolution
@@
This unique measurement strategy distinguishes BENQO from conventional VQAs, providing an alternative route for quantum optimization. To showcase that the benchmarking and scaling analysis is broadly applicable beyond standard VQA methods, BENQO is included in the comparative loss function studies in Section III.2.
@@ III Experimental Setup
    Problem Encoding: Random QUBO instances are used with the additional stipulation that the diagonal of the QUBO matrices contains only negative values. QUBO can be regarded as the standard formulation for combinatorial quantum optimization in the VQA context and provides a realistic benchmark (cf. Section III.1).
@@
    Quantum Subroutine: The loss function is derived from BENQO, as it exhibits a particularly well-behaved loss landscape compared to other established methods (cf. Section III.2).
@@
    Noise Modelling: Instead of relying on hardware-specific noise models, we introduce effective Gaussian noise directly to the loss function, ensuring a highly generalizable and systematic approach to model stochastic errors (cf. Section III.3) regardless of their origin.
@@
    Classical Subroutine: To eliminate biases arising from specific optimizer choices, we include a full suite of state-of-the-art methods in our experiments (cf. Section III.4).
@@
    Statistical Evaluation: By assigning a success index to each optimization run and repeating experiments $100$ times, we derive statistical measures of problem solvability, enabling a systematic scaling analysis (cf. Section III.5).
@@
Figure 3: Schematic overview of the experimental setup. (a) A random QUBO problem is sampled as input to the VQA pipeline (Section III.1). (b) The BENQO algorithm is simulated to obtain exact loss values for the given problem, chosen for its particularly well-behaved loss landscape (Section III.2). (c) The exact loss values are perturbed by additive Gaussian noise before being passed to a classical optimizer (Section III.3). (d) A suite of well-established classical optimization methods (Section III.4) attempts to find the optimal parameters minimizing the noisy loss function. (e) The results are then evaluated based on a predefined success criterion (Section III.5).
III.1 Random Problem Instances
@@
III.2 Well-Behaved Quantum Loss Functions
Trainability in VQA architectures is generally hindered by two key challenges: unfavorable loss landscape structures, which complicate optimization [13], and the barren plateau (BP) phenomenon, characterized by an exponential concentration of the loss function and its gradients [10]. While empirical studies on the former are computationally demanding due to the high dimensionality of the parameter space [49], BPs can be detected more efficiently by analyzing the sample variance of the quantum loss function [9, 50].
@@ III.2 Well-Behaved Quantum Loss Functions
To ensure that our scaling experiments represent a “best-case” scenario, we first conduct a preliminary study comparing three different VQA ansatzes, selecting the one with the most favorable loss landscape for the given problem class (cf. Section III.1). As a key indicator of trainability, we use the sample variance (14) of the loss functions, which provides insight into how well an optimizer can distinguish between different parameter settings.
Specifically, we compare a conventional QAOA implementation from Qiskit, a hardware-efficient two-local VQE ansatz, consisting of one rotational layer and one linear entangling layer, and the BENQO algorithm (cf. Section II.2). Each ansatz is parameterized such that the number of parameters is the same as the problem size $n$  [^1] This is a conservative assumption on parameter scaling, as larger problem instances typically require deeper circuits with more parameters, leading to higher-dimensional loss landscapes and potentially worse optimization conditions. Fixing this relationship therefore ensures a fair scalability comparison without biases from varying parameter growth., ensuring a fair comparison. For QAOA, this amounts to $p=n/2$ layers, and therefore requires $n$ to be even.
Fig. 4 presents the results of this variance analysis. While the empirical data may not conclusively confirm the presence of barren plateaus, the observable decay in loss variance still provides valuable comparative insights into algorithm trainability, as higher variance generally indicates better trainability. Among the three ansatz structures, BENQO consistently maintains the highest absolute variance in its loss values as the system size increases, suggesting the best parameter sensitivity. Given these results, BENQO is selected for the subsequent experiments, as it represents the most favorable case for classical optimization. For a more detailed comparison of the three algorithms that reinforces this conclusion, see Ref. [52].
@@
III.3 Scalable Gaussian Noise Model
@@ III.4 State-of-the-Art Optimizers
To ensure a fair comparison among all optimizers, we fix their available computational resources, using default settings from Qiskit. The initial parameters for each optimization run are sampled from a multivariate standard normal distribution centered at $\mathbf{0}\in\mathbb{R}^{n}$, a strategy shown to be effective in previous research [45, 46]. Each experiment consists of a single optimization run with a fixed optimizer (from Table 1), a specific noise level $\sigma$, and a random QUBO problem of size $n$. To account for statistical variation, we repeat each experiment $N=100$ times with different problem instances.
III.5 Measures of Solvability

### v2

@@ Abstract
With rapid advances in quantum hardware, a central question is whether quantum devices with or without full error correction can outperform classical computers on practically relevant problems. Variational Quantum Algorithms (VQAs) have gained significant attention as promising candidates in this pursuit, particularly for combinatorial optimization problems. While reports of their challenges and limitations continue to accumulate, many studies remain optimistic based on small-scale, idealized testing setups, leaving doubt about the scalability of VQAs for large-scale problems. We systematically investigate this scaling behavior by analyzing how classical optimizers minimize variational quantum loss functions for random QUBO instances in the presence of uncertainty, modeled as effective Gaussian noise. We find that the critical noise threshold for successful optimization decreases rapidly as system size grows. This decline exceeds what can be explained solely by shrinking loss variance, confirming deeper, fundamental limitations in the loss landscapes of VQAs beyond barren plateaus. Translating these thresholds into required measurement shots reveals that achieving sufficient precision in the evaluated loss values quickly becomes impractical, even for moderately-sized problems. Our findings demonstrate serious scalability challenges for VQAs in optimization stemming from mere uncertainty, indicating potential barriers to achieving practical quantum advantage with current hybrid approaches.
@@ I Introduction
While quantum hardware has advanced rapidly [1, 2, 3], scalable and fault-tolerant quantum computing may still be a long way off. John Preskill [4] coined the term NISQ (Noisy Intermediate-Scale Quantum) to describe a generation of quantum devices, which operate with limited resources and significant noise levels. An open question of the NISQ era is whether these devices can achieve a practically relevant quantum advantage. In this context, variational quantum algorithms (VQAs) have emerged as potentially promising candidates for realizing near-term quantum advantage in a wide range of applications [5, 6, 7]. One of the most anticipated use cases, particularly in industry, is combinatorial optimization, where current quantum approaches mainly rely on approximation algorithms and heuristics such as VQAs [8].
@@
VQAs are hybrid schemes that leverage quantum computers to evaluate a loss function while relying on classical optimization methods for loss minimization. They employ shallow, parameterized quantum circuits (ansatzes), making them well-suited to the resource constraints of NISQ devices. Yet, a growing body of research highlights significant challenges and limitations of VQAs: One major issue is the barren plateau phenomenon [9, 10], where the loss values and gradients vanish exponentially with system size in most of the proposed quantum ansatzes, making parameter optimization intractable. Conversely, structural properties leading to the provable absence of barren plateaus might enable efficient classical simulations [11, 12]. Additionally, the classical optimization problem within VQAs has been proven to be NP-hard [13], with loss landscapes plagued by numerous suboptimal local minima [14], further complicating the convergence to high-quality solutions. These challenges are amplified by noise, which not only leads to unfavorable resource scaling and unpredictable behavior but may also eliminate any potential quantum advantage altogether [15, 16, 17, 18, 19, 20, 21]. Crucially, these limitations persist regardless of whether the effective noise comes from hardware imperfections or finite sampling errors inherent to quantum measurement. Consequently, despite ongoing efforts and widespread optimism, no practically meaningful quantum advantage has been demonstrated for VQAs on near-term devices to date.
@@
Thus, the critical question remains: can variational quantum optimization become practical in the presence of unavoidable stochastic errors? While some studies directly address this question (see Section II), most research focuses on developing new heuristic approaches, typically tested on small-scale and idealized setups. Such benchmarks often yield promising insights but rarely capture how well algorithms scale under realistic noise conditions and increasing problem sizes, potentially painting an overly optimistic picture of their viability. More importantly, many studies overlook the role of the classical optimization process – the central computational bottleneck of VQAs – which may ultimately determine whether a problem remains solvable as system size grows. Closing this gap is essential for distinguishing genuinely promising advancements from appealing but unproven claims in quantum optimization and related disciplines.
@@
This work investigates the viability of VQAs for solving classical optimization problems in the Quadratic Unconstrained Binary Optimization (QUBO) formulation, with a specific focus on the classical optimization routine as a key driver of performance. We systematically analyze how state-of-the-art classical optimizers handle variational quantum loss functions under increasing levels of Gaussian noise, providing an empirical assessment of practical scalability limits in variational quantum optimization. The primary contributions of this study are:
@@
    Performance Evaluation Framework: We propose a systematic methodology to assess the scaling behavior of hybrid variational quantum optimization methods under stochastic noise, providing a robust foundation for understanding their potential and limitations, even in fault-tolerance.
@@
    Scaling Analysis of Classical Optimizers: This study presents, to the best of our knowledge, the first in-depth scaling analysis of classical optimizer performance on noisy quantum loss functions for random QUBO problem instances. By examining two essential factors – noise level ($\sigma$) and system size ($n$) – we identify a clear threshold $\sigma^{*}(n)$, beyond which reliable optimization becomes exponentially challenging. Via extrapolation to regions not (yet) accessible numerically we address the feasibility of these algorithms for larger-scale problems.
@@
    Insights on Practical Quantum Advantage: We quantify how uncertainty in the loss values, particularly arising from finite measurement sampling, can significantly impact the practical scalability of VQAs. These results highlight potential barriers to achieving quantum advantage in variational quantum optimization under realistic noise scenarios.
Our empirical findings offer clear benchmarks and valuable insights into the practical limits of VQAs under noise. By fostering a more realistic understanding of the potential of quantum optimization on near-term devices and beyond, these insights help to guide future research towards more promising algorithms and applications in quantum computing.
II Related Work
The trainability of VQAs has been extensively explored in the literature. Several key obstacles impairing successful parameter optimization have been identified, including reachability deficits [22], inherent NP-hardness of optimization [13], complex loss landscapes with numerous suboptimal local minima [14], and barren plateaus [9, 10], i.e. the phenomenon of exponentially vanishing loss values and gradients. Specifically, recent work indicates that all sources of barren plateaus, such as high ansatz expressibility [9], global observables [23], or hardware-induced noise [16], are fundamentally connected to the scaling behavior of the dynamical Lie algebra (DLA) associated with the chosen ansatz [23, 24, 25, 26]. Our study explicitly circumvents these issues by considering only ansatz circuits characterized by only polynomially growing DLA dimensions along with strictly local observables, ensuring that the underlying quantum loss functions are provably barren-plateau-free and remain “trainable” in theory.
A significant body of research has also addressed the detrimental effects of hardware noise on VQA performance. Numerous studies have demonstrated that even minimal noise levels can rapidly propagate through quantum circuits, resulting in output distribution that become effectively simulable classically [19, 27, 17, 28, 29, 20, 21], thereby negating potential quantum advantage. In contrast, our analysis assumes a fault-tolerant scenario, simulating uncertainty exclusively via Gaussian noise applied as a post-processing step to exact quantum loss function evaluations. This setup effectively models an idealized, fully error-corrected scenario with finite sampling noise as the dominating source of uncertainty. This setting not only circumvents previously identified noise-induced limitations but also aligns more closely with realistic conditions under which quantum advantage for (approximate) optimization can be expected to first emerge.
Given the empirical and experimental nature of our study and its focus on the classical optimization routine, our findings naturally compare to existing numerical assessments of classical optimizer performance within VQAs [30, 31, 32, 33, 34, 35, 36, 37, 38]. However, as already noted in Ref. [39], these studies often evaluate performance under fixed conditions using static setups primarily focusing on comparing optimizer effectiveness. Consequently, they offer limited insights into scalability constraints and the quantitative limits of classical optimizer performance in hybrid quantum algorithms with respect to noise thresholds or system size.
Most closely aligned with our work in terms of scope and metrics is the more recent study by Scriva et. al. [18], which employs similar performance metrics and comes to similar conclusions. However, methodological differences exist: Scriva et. al. [18] focus exclusively on shot noise and evaluate iteration complexity relative to fixed success probabilities. Conversely, our approach systematically investigates the effects of varying levels of Gaussian noise, allowing for a broader and more generalizable analysis of the precision requirements for successful optimization. A complementary theoretical investigation into the convergence properties of classical optimizers within a similar context is provided by Kungurtsev et. al. [40].
III Theoretical Background
@@
III.1 Quadratic Unconstrained Binary Optimization Problems
@@
III.2 Variational Quantum Optimization
@@
Fig. 1 illustrates representative ansatz circuits for both algorithm classes, which are also used in the experiments presented in Section IV.2. For VQE, a typical example is a hardware-efficient ansatz, consisting of alternating layers of single-qubit rotations and entangling operations, followed by a global measurement. QAOA, in contrast, follows a problem-inspired approach, implementing a trotterized version of adiabatic quantum optimization [46]. It consists of alternating applications of a shift operator ${\mathbf{U}_{P}(\gamma)=e^{-i\gamma\mathbf{H}_{P}}}$ based on the problem Hamiltonian $\mathbf{H}_{P}$, and a mixing operator ${\mathbf{U}_{M}(\beta)=e^{-i\beta\mathbf{H}_{M}}}$, where the default choice is a sum of Pauli-$\mathbf{X}$ matrices on each qubit, i.e., ${\mathbf{H}_{M}=\sum_{i=1}^{n}\mathbf{X}_{i}}$. The resulting QAOA ansatz is composed of $p$ layers, parameterized by $2p$ angles, with the unitary evolution
@@
This unique measurement strategy distinguishes BENQO from conventional VQAs, providing an alternative route for quantum optimization. To showcase that the benchmarking and scaling analysis is broadly applicable beyond standard VQA methods, BENQO is included in the comparative loss function studies in Section IV.2.
@@ IV Experimental Setup
    Problem Encoding: Random QUBO instances are used with the additional stipulation that the diagonal of the QUBO matrices contains only negative values. QUBO can be regarded as the standard formulation for combinatorial quantum optimization in the VQA context and provides a realistic benchmark (cf. Section IV.1).
@@
    Quantum Subroutine: The loss function is derived from BENQO, as it exhibits a particularly well-behaved loss landscape compared to other established methods (cf. Section IV.2).
@@
    Noise Modelling: Instead of relying on hardware-specific noise models, we introduce effective Gaussian noise directly to the loss function, ensuring a highly generalizable and systematic approach to model stochastic errors (cf. Section IV.3) regardless of their origin.
@@
    Classical Subroutine: To eliminate biases arising from specific optimizer choices, we include a full suite of state-of-the-art methods in our experiments (cf. Section IV.4).
@@
    Statistical Evaluation: By assigning a success index to each optimization run and repeating experiments $100$ times, we derive statistical measures of problem solvability, enabling a systematic scaling analysis (cf. Section IV.5).
@@
Figure 3: Schematic overview of the experimental setup. (a) A random QUBO problem is sampled as input to the VQA pipeline (Section IV.1). (b) The BENQO algorithm is simulated to obtain exact loss values for the given problem, chosen for its particularly well-behaved loss landscape (Section IV.2). (c) The exact loss values are perturbed by additive Gaussian noise before being passed to a classical optimizer (Section IV.3). (d) A suite of well-established classical optimization methods (Section IV.4) attempts to find the optimal parameters minimizing the noisy loss function. (e) The results are then evaluated based on a predefined success criterion (Section IV.5).
IV.1 Random Problem Instances
@@
IV.2 Well-Behaved Quantum Loss Functions
Trainability in VQA architectures is generally hindered by two key challenges: the barren plateau (BP) phenomenon, characterized by an exponential concentration of the loss function and its gradients [10], and other unfavorable loss landscape structures which complicate optimization [13, 14]. While empirical studies on the latter are computationally demanding due to the high dimensionality of the parameter space [68], BPs can be detected more efficiently by analyzing the sample variance of the quantum loss function [9, 24].
@@ IV.2 Well-Behaved Quantum Loss Functions
In Fig. 4, we illustrate the scaling of variances across three different VQA ansatzes: a conventional QAOA implementation from Qiskit, a hardware-efficient two-local VQE ansatz, consisting of one rotational layer and one linear entangling layer, and the BENQO algorithm (cf. Section III.2). Each ansatz is parameterized such that the number of parameters matches the problem size $n$  [^1] This is a conservative assumption on parameter scaling, as larger problem instances typically require deeper circuits with more parameters, leading to higher-dimensional loss landscapes and potentially better optimization conditions [136]. Fixing this relationship therefore ensures a fair scalability comparison without biases from varying parameter growth., ensuring a fair and unbiased comparison. Specifically, for the QAOA ansatz this amounts to $p=n/2$ layers, and therefore requires $n$ to be even.
Based on a characterization of their dynamical Lie algebra (DLA) structures, we know neither the Two Local ansatz used here nor the problem-dependent BENQO ansatz exhibit exponentially scaling DLA dimensions [70, 71]. Conversely, QAOA when applied to the Ising cost function defined in Eq. 2 is known to have an exponentially scaling DLA dimension and therefore exhibits a barren plateau [24, 70]. While various mitigation strategies, such as smart parameter initialization methods (discussed in detail in Section VII.4), can substantially improve the trainability of such ansatzes, this study explicitly concentrates on ansatzes that inherently avoid barren plateaus. Among the two remaining ones tested, BENQO consistently maintains the highest absolute variance in its loss values with increasing system size, suggesting the best parameter sensitivity. Therefore, BENQO is selected for the subsequent experiments as a representative scenario for a favorable case for classical optimization.
@@
IV.3 Scalable Gaussian Noise Model
@@ IV.4 State-of-the-Art Optimizers
To ensure a fair comparison among all optimizers, we fix their available computational resources, using default settings from Qiskit. The initial parameters for each optimization run are sampled from a multivariate standard normal distribution centered at $\mathbf{0}\in\mathbb{R}^{n}$, a strategy shown to be effective in previous research [86, 64, 65]. Each experiment consists of a single optimization run with a fixed optimizer (from Table 1), a specific noise level $\sigma$, and a random QUBO problem of size $n$. To account for statistical variation, we repeat each experiment $N=100$ times with different problem instances.
IV.5 Measures of Solvability